# PlanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

### v1

@@ Abstract
We introduce PlanQA, a diagnostic benchmark for evaluating geometric and spatial reasoning in large-language models (LLMs). PlanQA is grounded in structured representations of indoor scenes, such as kitchens, living rooms, and bedrooms, encoded in a symbolic format (e.g., JSON, XML layouts). The benchmark includes diverse question types that test not only metric and topological reasoning (e.g., distance, visibility, shortest paths) but also interior design constraints such as affordance, clearance, balance, and usability. Our results across a variety of frontier open-source and commercial LLMs show that while models may succeed in shallow queries, they often fail to simulate physical constraints, preserve spatial coherence, or generalize under layout perturbation. PlanQA uncovers a clear blind spot in today’s LLMs: they don’t consistently reason about real-world layouts. We hope that this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
@@ 1 Introduction
Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over physically grounded environments remains poorly understood. In numerous real-world applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language. Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
We introduce PlanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floor plans represented in structured text-based formats. Each instance consists of a JSON-encoded layout paired with natural language questions that require the model to compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints. PlanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
Although LLMs can increasingly be used in tool-assisted pipelines, for example to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities. PlanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding.
@@
  • We introduce a dataset of 1,800 structured 2D layouts, 600 each from kitchens, living rooms, and bedrooms, represented in JSON and paired with spatial reasoning questions.
  • We provide a diverse suite of 18,000 questions spanning geometric, functional, and design-aware reasoning, with one question per layout across multiple categories.
@@
  • We conduct a comparative analysis of 11 LLMs, including GPT-4.1 OpenAI et al. (2024) and open models such as DeepSeek-R1 DeepSeek-AI (2025) and Llama-3.3-70B-Instruct Grattafiori et al. (2024), revealing consistent failure patterns in spatial inference from symbolic input.
PlanQA provides a diagnostic substrate for measuring spatial competence in language models, with applications in layout validation, agent planning, and symbolic grounding.
@@ 2 Related Work
Prior benchmarks have explored spatial reasoning across vision and language domains. CLEVR Johnson et al. (2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relationships. In real-world settings, SpatialSense Yang et al. (2019) focuses on recognizing spatial relations in images through adversarially mined examples. Benchmarks like BabyAI Chevalier-Boisvert et al. (2019), ALFRED Shridhar et al. (2020), and Room-to-Room (R2R) Anderson et al. (2018) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments. Recent datasets such as ScanQA Azuma et al. (2022) and 3DSRBench Ma et al. (2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
@@
Advancements in generative models have also contributed to spatial reasoning tasks. LayoutGPT Feng et al. (2023) leverages large language models for compositional visual planning and layout generation, while Holodeck Yang et al. (2024) enables language-guided generation of 3D embodied AI environments. Similarly, AnyHome Fu et al. (2024) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts. Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but suffers from poor object placement, often yielding physically implausible or random layouts due to non-convergent simulated annealing. LayoutVLM Sun et al. (2024) and FirePlace Huang et al. (2025) improve layout generation via optimization and constraint solving, respectively. But they assess spatial reasoning through output realism, not direct inference. In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models Yamada et al. (2024), which assesses the spatial reasoning capabilities of LLMs through structured tasks. Additionally, benchmarks such as BALROG Paglieri et al. (2025) test agentic reasoning in game environments, further exploring the spatial and decision-making abilities of language and vision-language models. While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning in structured layouts, providing fine-grained and task-specific insights into models’ spatial competence. Recent 3D-LLM surveys such as Ma et al. (2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning. PlanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
PlanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts. Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, PlanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
@@ 3.1 Synthetic Layout Generation
Our initial aim was to utilize publicly available real-world floorplan datasets. However, a comprehensive review revealed critical limitations. Prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) were withdrawn due to unresolved copyright claims, rendering them unusable. Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2020), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse. Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations. RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al. (2020), while large, appears procedurally generated albeit with realtor supervision and imposes restrictions on redistribution and derived works. Given these legal, practical, and methodological constraints, we found synthetic data generation to be the most viable alternative.
We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Team et al., 2025). While employing an LLM for both data generation and evaluation may raise concerns of circularity, our evaluation pipeline is entirely independent: the LLM is used solely for data generation, and evaluation is performed via deterministic spatial reasoning tasks.
@@
The generation process comprises two stages. First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to circulation, symmetry, and zoning. These constraints are encoded directly in the LLM prompt. Second, each room is furnished according to style-specific rules, also defined in structured prompts, to encourage both visual realism and functional plausibility. Approximately one-third of candidate layouts are filtered out by automated spatial validity checks, which remove configurations with inaccessible furniture or implausible adjacencies. This curation pipeline yields a diverse and geometrically coherent dataset. Full prompts, generation templates, and validation scripts are provided in the supplemental materials. Representative examples of layout types are shown in Fig. 1. A breakdown of room types and layout geometries is provided in Table 1.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Representative layouts from the PlanQA dataset, spanning three room types (kitchen, living room, bedroom) and three geometric configurations (rectangular, L-shaped, open). Layouts are procedurally generated using explicit spatial constraints and validated for functional feasibility.
@@
Table 1: Layout distribution in PlanQA by room type, style, and geometry, with room-wise totals.
@@
Kitchen L-Shaped (incl. counter) 208 36 105 600
U-Shaped / G-Shaped 42 128 2
@@
Wall / Galley (compact layouts) 17 10 17
@@
Bedroom Bedroom with Workstation 98 35 28 600
Traditional Bedroom 114 59 57
Efficient Small Bedroom 65 3 8
Welcoming Guest Bedroom 75 30 28
@@
3.2 Question Taxonomy, Prompting, and Cognitive Scope
@@
PlanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout. Questions span a range of geometric, topological, and functional types, including numeric computations (distances, areas), spatial feasibility (object fit, placement), occlusion and visibility, and environment-level assessments such as missing elements or requirement violations. Examples include queries like whether a 2 m $\times$ 3 m table can be placed without blocking passage, whether the television is visible from a particular seat, or which essential appliances are absent from a proposed kitchen. Some questions require local metric reasoning; others test whether a model can simulate constraints that would arise in physical occupancy or usability. A categorized list of question types is shown in Table 2.
Each question is generated by instantiating a parameterized template against the specific geometry of a layout. These templates are natural language prompts that fill in object names, measurements, and task-specific context. Prompts are issued in zero-shot settings with no few-shot scaffolding or role-based instruction, but include consistent structural cues to support stepwise inference. Each prompt ends with the string ”Reason step by step to solve the problem.” followed by “Final Answer:”, enabling unambiguous output extraction. For instance, a placement query might read: “Given the room layout below, compute the area of the largest rectangle that can be placed in the kitchen without overlapping any object. Reason step by step to solve the problem. Final Answer: [value in square meters].”
@@
Table 2: PlanQA question taxonomy. The example question shown is an instantiation of the template used to generate all questions of that type. Each task is labeled with a format code, cognitive skill category (see Tables 4 and 4), and the room types in which it is applied: K = kitchen, L = living room, B = bedroom.
Category Example Question Format Target Rooms
Distance What is the distance between the refrigerator and the stove? N M K, L, B
Area What is the total sitting area? N M K, L, B
Free Space What is the unoccupied floor area in the room? N T K, L, B
View Angle What is the angle between the TV’s facing direction and the viewer’s line of sight from the armchair? N V K, L, B
@@
Repositioning How far can the ottoman be moved to the left direction until it touches another object or the wall? N G K, L, B
Max Box What is the area of the largest axis-aligned rectangle that fits in the free space? N G K, L, B
@@
Fit/Placement Can a 2m × 3m desk table fit in the room without overlap? B G K, L, B
Shortest Path What is the valid shortest path from the stove to the exit with clearance? S G, R K, L, B
Missing Object What is a possible placement for the missing sink? S R, T K, L, B
Obstruction What are all the objects that intersect the straight line from the fireplace to the door? L V K, L, B
Table 3: Answer format and scoring codes.
Code Format Scoring
N Scalar Rel. error $\leq$ 2%
B Boolean Exact match
S Sequence Fréchet $\leq$ 0.6 m
L List Set equality
C Coord. Spatial predicate
Table 4: Cognitive skill clusters.
Code Cognitive Target
M Metric computation
T Topology
G Geometric planning
V Visual logic / occlusion
R Rule-based validation
Structured layouts are flattened into lists of labeled objects, each defined by an axis-aligned bounding box. Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents are unique and stable across prompt construction and answer evaluation. All coordinates are expressed in meters, using a bottom-left origin and $(x,y)$ ordering. There is no hierarchical or typed schema beyond the flat object list, and no implicit conventions are assumed. A single system prompt is prepended to every interaction, specifying general and room-specific constraints: “You are a spatial assistant. Assume that all windows have zero thickness.” For layouts of type living room or bedroom, the prompt further clarifies that rugs may be overlapped by other objects, allowing for more realistic spatial arrangements involving soft surfaces.
Some question types are instantiated deterministically given the layout geometry (e.g., exact distances between centroids), while others are drawn from constrained samplers over feasible object pairs, free-space regions, or task-specific heuristics. The prompt templates themselves remain fixed across layouts and models. This uniformity allows us to isolate spatial reasoning under consistent linguistic and structural conditions.
These cognitive categories correspond to distinct reasoning faculties. Metric tasks (M) require explicit numerical computations over distances, areas, or positions, often grounded in real-valued geometric attributes. Topological tasks (T) depend on spatial relationships like adjacency or containment that are invariant under metric transformations. Geometric planning (G) involves multistep spatial inference over object configurations, including feasibility, collision avoidance, or spatial optimization (e.g., placing a large box without overlap). Visual logic (V) tasks such as visibility and angle estimation rely on implicit occlusion reasoning or directional projection. Rule-based tasks (R) involve symbolic constraints, such as verifying whether a layout satisfies clearance or adjacency requirements or whether essential objects are missing. These clusters are intended to reflect the diversity of tasks while enabling diagnostic grouping of error types. They also reflect a spectrum of abstraction, fromm low-level metric properties to high-level design affordances. While these clusters are not mutually exclusive, they provide a useful abstraction for analyzing model behavior and diagnosing failures. For example, we hypothesize that questions involving geometric planning and occlusion (G, V) will be disproportionately challenging for models without grounded spatial priors. This hypothesis is revisited in our results analysis (§4). The formatrmat and target codes are defined in Tables 4 and 4, and used throughout Table 2.

### v2

# FloorplanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

@@ Abstract
We introduce FloorplanQA, a diagnostic benchmark for evaluating spatial reasoning in large-language models (LLMs). FloorplanQA is grounded in structured representations of indoor scenes, such as (e.g., kitchens, living rooms, bedrooms, bathrooms, and others), encoded symbolically in JSON or XML layouts. The benchmark covers core spatial tasks, including distance measurement, visibility, path finding, and object placement within constrained spaces. Our results across a variety of frontier open-source and commercial LLMs reveal that while models may succeed in shallow queries, they often fail to respect physical constraints, preserve spatial coherence, though they remain mostly robust to small spatial perturbations. FloorplanQA uncovers a blind spot in today’s LLMs: inconsistent reasoning about indoor layouts. We hope this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
@@ 1 Introduction
Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over plausible, physically feasible environments such as indoor layouts remains poorly understood. In numerous practical applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language. Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
We introduce FloorplanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floorplans represented in structured text-based formats. Each instance consists of a JSON-encoded layout paired with natural language questions that require the model to compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints. FloorplanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
Although LLMs can increasingly be used in tool-assisted pipelines, for example to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities. FloorplanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding, in order to measure their unassisted capabilities. This baseline is important because even in tool-rich systems, models benefit from some unaided spatial ability to anticipate outputs and avoid trivial errors.
@@
  • We introduce a dataset of 2,000 structured 2D layouts, including 600 each from synthetically generated kitchens, living rooms, and bedrooms, plus 200 layouts sourced from the Habitat Synthetic Scenes Dataset (HSSD) (Khanna et al., 2023), providing a realism check. All are represented in JSON and paired with spatial reasoning questions.
  • We provide a diverse suite of 16,000 spatial reasoning questions, eight questions per layout, covering geometric relations, placement feasibility, spatial occupancy, and navigation.
@@
  • We conduct a comparative analysis of 15 LLMs, including 7 reasoning-focused models, as well as 8 standard models, revealing consistent failure patterns in spatial inference from symbolic input.
FloorplanQA provides a benchmark of layouts, questions, and evaluation metrics for assessing spatial reasoning in language models, focusing on symbolic floorplans that integrate geometry and semantics in ways that mirror real architectural abstractions.
@@ 2 Related Work
Prior benchmarks have explored spatial reasoning across vision and language domains. CLEVR Johnson et al. (2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relations. In real-world settings, SpatialSense Yang et al. (2019) focuses on recognizing spatial relations in images through adversarially mined examples. Benchmarks like BabyAI Chevalier-Boisvert et al. (2019), ALFRED Shridhar et al. (2020), and Room-to-Room (R2R) Anderson et al. (2018) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments. Recent datasets such as ScanQA Azuma et al. (2022) and 3DSRBench Ma et al. (2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
@@
Advancements in generative models have also contributed to spatial reasoning tasks. LayoutGPT Feng et al. (2023) leverages large language models for compositional visual planning and layout generation, while Holodeck Yang et al. (2024) enables language-guided generation of 3D embodied AI environments. Similarly, AnyHome Fu et al. (2024) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts. Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but often produces implausible object placement due to non-convergent simulated annealing. LayoutVLM Sun et al. (2024) and FirePlace Huang et al. (2025) improve layout generation via optimization and constraint solving, respectively. But they assess output realism, not the model’s ability to infer constraints directly. In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models Yamada et al. (2024), which assesses the spatial reasoning capabilities of LLMs through structured tasks. Additionally, benchmarks such as BALROG Paglieri et al. (2025) test agentic reasoning in game environments, further exploring the spatial and decision-making abilities of language and vision-language models. While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning in structured layouts, providing fine-grained and task-specific insights into models’ spatial competence. Recent 3D-LLM surveys such as (Ma et al., 2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning. FloorplanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
FloorplanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts. Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, FloorplanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
@@ 3.1 Synthetic Layout Generation
Our initial aim was to use publicly available real-world floorplan datasets. However, a comprehensive review revealed significant limitations, as several prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) are not accessible due to unresolved copyright claims. Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2020), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse. Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations. RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al. (2020), while large, is procedurally generated with realtor supervision and imposes restrictions on reuse. Given these legal, practical, and methodological constraints, we found synthetic data generation to be the most viable alternative.
We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Google Gemini Robotics Team, 2025). Although our evaluation includes multiple LLMs (including Gemini variants), there is no circularity: the data-generation step is one-off and separate from evaluation. For evaluation, models solely produce answers to the benchmark questions, and correctness is computed by deterministic geometric routines against ground-truth solutions derived from the layouts, so no model output is used to generate data or to grade itself.
@@
The generation process comprises two stages. First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to circulation, symmetry, and zoning. These constraints are encoded directly in the LLM prompt. Second, each room is furnished according to style-specific guidelines (e.g., a bedroom must contain a bed and storage), also defined in structured prompts, to encourage both visual realism and functional plausibility. Approximately one-third of candidate layouts are filtered out by a rule-based spatial validity filter that enforces basic clearance and accessibility constraints. The checks remove scenes with inaccessible furniture and implausible adjacencies, such as sofas blocking doors or a refrigerator overlapping a table; see Appendix B.3 for the full set of cases. Full prompts, generation templates, and validation scripts are provided in the Supplementary Material.
3.2 Layout Extraction from HSSD Dataset
To complement the synthetically generated layouts, we further incorporated 200 layouts extracted from the Habitat Synthetic Scenes Dataset (HSSD-200). HSSD provides 211 high-quality, human-authored 3D scenes designed with the Floorplanner interface and populated with 18,656 objects across 466 semantic categories. Unlike purely procedural datasets, HSSD offers fine-grained semantics, 3D assets, and close correspondence to real interiors, making it an effective proxy for real-world interior layouts.
For our purposes, we project each 3D scene to a 2D floorplan and retain only select structural and furniture elements. Decorative or auxiliary objects (e.g., vases, plants, cushions, artworks, posters, bottles, shoes, candles) are removed to reduce clutter; see Appendix C for the full details. We then use an $\alpha$-convex hull (Asaeedi et al., 2014; Edelsbrunner & Mücke, 1994) to smooth object boundaries, yielding polygonal layouts that are not restricted to axis-aligned rectangles. This step is necessary because raw HSSD projections often produce overly dense polygons, with redundant vertices along straight or nearly straight segments; applying an $\alpha$-hull reduces spurious complexity while preserving concavity, which avoids unnecessary token overhead in downstream LLM processing. This ensures compatibility with our synthetic layouts, while maintaining the richer geometric variety of HSSD.
3.3 Unified Dataset
Together, these two sources yield a dataset of 2,000 layouts: 1,800 synthetically generated via Gemini 2.5 Pro and 200 extracted from HSSD. The two subsets share a unified polygonal representation, enabling consistent downstream processing. Figure 1 shows examples from both sources while Table 1 summarizes room, style, and geometry distributions across the synthetic subset with the figure next to it providing a breakdown by room type.
@@ 3.3 Unified Dataset
Table 1: Layout distribution in FloorplanQA by room type, style, and geometry, with room-wise totals.
@@
Kitchen L-Shaped 208 36 105 600
U/G-Shaped 42 128 2
@@
Wall / Galley 17 10 17
@@
Bedroom With Workstation 98 35 28 600
Traditional 114 59 57
Efficient Small 65 3 8
Welcoming Guest 75 30 28
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Representative layouts from FloorplanQA. Generated: kitchen, living room, bedroom. HSSD: last image. Generated objects are axis-aligned boxes; HSSD uses arbitrary polygons.
3.4 Question Taxonomy and Prompting
@@
FloorplanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout. Questions span a range of topological and functional types, including numeric computations (distances, areas), spatial feasibility (object placement), visibility, and requirement violations. Some questions require fine-grained metric reasoning, others test whether a model can respect physical constraints. A categorized list of question types is shown in Table 2.
Each question is generated by filling a parameterized template with layout-specific variables such as object names, measurements, and task-specific contextual information (e.g. units for distance, clearance for paths, or which object-types should not occlude visibility). Prompts are issued in zero-shot settings, without few-shot examples or role-based instructions. Instead, we enforce simple structural markers—such as a required checklist and a final-answer line—to encourage stepwise reasoning.
To ensure verifiable outputs, each prompt specifies a response schema consisting of a brief structured justification and a final answer line. For example, a distance query is phrased as:
Prompt: Distance Query If the format, object names, or required inputs are missing, invalid, or inconsistent, the model must return:  *Final answer*: ERROR. Otherwise, responses must follow the scheme, for example: Response Schema This structure invokes step-by-step reasoning and each question ends with *Final Answer*: <answer>, enabling robust extraction even when APIs lack native structured-output support.
@@
Table 2: FloorplanQA question taxonomy. The example question shown is an instantiation of the template used to generate all questions of that type. Each task is labeled with a format code: N (scalar), B (boolean), S (sequence), and L (list), and a question reasoning category.
Type Example Question Format Category
Distance Calculate the Euclidean distance in meters between the centroids of the fridge and the stove N Metric
Free Space Calculate the total non-occupied floor area in square meters N Topology
View Angle Compute the smallest absolute angle in degrees between the vector from the centroid of the sofa to the centroid of the TV and the global north vector (0, 1) N Metric
@@
Repositioning Calculate how far the ottoman be moved in the left direction until it touches another object or the wall N Dynamic
Max Box Calculate the area in square meters (m²) of the largest rectangle that can fit inside the room N Topology
@@
Placement Check if a 2m × 3m desk table can fit fully inside the room without overlaps B Topology
Shortest Path Determine the shortest valid path that maintains a clearance of 15 cm from all other objects, starting from centroid of the stove and ending at the centroid of door S Dynamic
Visibility Find all objects that intersect the vector from the centroid of the window to the centroid of the fireplace L Topology
Layouts are represented in a structured JSON format. Each entry contains a layout_id, the room_type, and explicit geometric descriptions. The room_boundary is stored as a closed polygon, while walls are represented separately as a list of line segments. Openings such as windows and doors are included in a dedicated openings field rather than flattened into the object list. All furnishings and functional elements (e.g., bed, sofa, table) are stored in the objects list, with each object defined by a labeled polygon. In the synthetic data, these polygons are axis-aligned bounding boxes (four points), whereas in HSSD they can exhibit arbitrary shapes and orientations. Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents remain unique and stable across prompt construction and answer evaluation. Coordinates are expressed in meters in a right-handed 2D Cartesian frame (+x+x to the right, +y+y upward). The global origin $(0,0)$ is not fixed to the layout’s lower-left corner and may vary across layouts. The prompt used to generate examples according to this schema is in Appendix H. We group these questions into three reasoning categories. Metric tasks require explicit numerical computation, such as calculating centroids, measuring distances between objects, or evaluating the angle between an inter-object vector and a reference axis. Topology category involves geometric and relational reasoning, including checking placement feasibility, computing free space, or identifying whether an object blocks the direct line of sight between two others. Dynamic category addresses layout-changing procedures, such as repositioning an object until contact with a boundary or another object, or computing a valid collision-free path between two objects. These categories are intended to capture the core modes of spatial reasoning in FloorplanQA, ranging from low-level geometric calculation to higher-level relational and procedural inference. While not strictly disjoint, they provide a diagnostic framework for analyzing model behavior and diagnosing failure modes. In addition to categorizing by reasoning type, each task is also associated with an answer format code that specifies the expected output structure and the corresponding scoring rule. Scalar outputs (N) are scored by relative error with a default tolerance of 2%; for complex area-computation tasks (e.g., Free Space), the tolerance is relaxed to 5%. Sequence outputs (S) are evaluated with a Fréchet threshold of 0.6 m, approximating minimal human clearance, and must be valid (collision-free, no overlaps). List outputs (L) are evaluated by set equality. Together, the categories and format codes define the taxonomy summarized in Table 2, which reports task coverage and examples for each case.